Short Report – QA/QC Meeting Feb 2018

Date: 2018-02-22

Participants / Excused


Talking Points


  1. Please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.
  2. From now on, Short Reports will be generated for all QA/QC Meetings.
  3. Short statement about site from SRP and Tech-SRP needed (max. 2 sentences)
  4. Starting somewhen later this year, weekly checks will be done online via checklists instead of the server Excel sheet
  5. Please be patient with QV changes, priority conflicts w/ ICOS.
  6. Automatic fluxes will be activated again (goal is March 2018).
  7. QuickViews Wishlist: what do you want for the QuickViews? Please send your wishes to LH.
  8. CH-FRU: what should LH check? LH checks meteo and fluxes.


  1. Not enough space on group drive for RDS_calculations, problem when converting raw binaries from one complete year. Group drive is a temporary (very short-term storage) folder.
    Disk space snapshots of flux folders, 16 Feb 2018:
    RDS_calculations: CH-LAE: 120 GB // CH-OE2: 57 GB // CH-CHA: 0.4 GB // all other sites: less than 0.4 GB // Total: 200 GB
    LH checks if possible to get own volume for temporary flux calcs.
  2. Folder structure on grassland server, GL-RAW needs cleaning (see Images). Not high priority, but should be cleaned at some point.
  3. Flux meteo screening tool maybe not fully compatible with flux processing tools, needs check. Topic moved to next meeting.
  4. The subfolders for CH-FRU on GL-RAW seem to be wrongly labelled as R3_LI72, it is a IRGA75 (see Images). Not high priority, but should be cleaned at some point.





QuickViews Check

LH checks if QVs were succesfully generated, at least once a month, for all sites.

Problems for QVs generation, 21 Feb 2018

  • Windows 10 on VPC updates, shuts down the VPS and thus logs out the QV user, QVs generation stops
  • QVs still try to access the old csv data files instead of the new Loggernet data files (at CH-AWS?)

Generation Checklist, 21 Feb 2018

CH-AWS Problems, Meteo script ran OK yesterday, but not today, and data stops on 1 Feb, probably b/c now Loggernet data files?
CH-CHA Problems, Meteoscript ran OK yesterday, but not today
CH-DAV Problems, Meteoscript ran OK yesterday, but not today
CH-FRU Problems, Meteoscript ran OK yesterday, but not today; Webcam pic is out-of-date
CH-LAE Problems, Meteoscript ran OK yesterday, but not today
CH-OE2 Problems, Meteoscript ran OK yesterday, but not today

Site Reports


SRP (LH): Fluxes look OK, data availability is currently nearly perfect. At some point we should investigate how the adjusting measurement height affects fluxes and the footprint.
Tech-SRP (IB): All instruments work fine, only the mast is not moving as it should, therefore the mast height is not adapted to the snowheight. Adapted it manually on the 31st of January 2018, therefore it stayed adapted until 15th of February since there was no snowfall during that time.


SRP (KF): Chamau fluxes for CO2 seem ok, daily cycle and CO2 uptake during daytime already in January. We have the Laser working again, interrupted once in February when software was running but data transfer to the embedded system, N2O fluxes close to zero. Rain gauge was failing in January, its connection to the logger was broken and fixed by PL.
Tech-SRP (PL): QCL data loss because serial feed was interrupt. Flux data loss for 1.5 days because of embedded computer system software updates. Rain gauge data loss because of a bitten cable (Mice). GF4 and GF5 soil sensors are partially showing bad data (sensors at end of life?)


SRP (LH): Data gap of several days for N2O and CH4 EC fluxes, the QCL had some problems. IRGA fluxes OK, for H2O we can see a clear diel cycle.
Tech-SRP (PM): BF-5 diffuse PAR sensor was installed at the site. Several soil flux chamber problems, often due to heavy snowfall. QCL problems. Continuous work on the storage flux profile.


SRP (LH): CH-FRU uses the LI-7500, winter CO2 fluxes with large spikes. H2O fluxes seem off around 7 Jan, but look better afterwards. Some data gaps, probably due to the AGC being too low.
Tech-SRP (PL): TS_GF1_0x02_1 has broken. New TS probe (TS_GF1_***_2) is wired wrong and therefore not measuring correctly.


SRP (EPL): LAE: The fluxes could not be calculated due to problems on RDS, which we are working on fixing. The LI-7500 was permanently taken down on December 12, 2017. The LI-7200 was taken down for calibration on December 12, 2017 and reinstalled on December 14, 2017. The subcanopy system (LAS) was reinstalled on 05/02/2018 and the data is automatically sent to ETH. When reinstalled the time was off by about 4 minutes.
Tech-SRP (TB): Last few task as mentioned from EPL and according to the fieldbook. Besides that I couldn’t install the newly repaired LI7500 since it was not upstarting in the cold [and] slowly degrading of our «old» equipment as mentioned also in the techmeeting.


SRP (CE): There is continuous flux data for this year so far, however noisy sonic anemometer data (as already experienced in the two previous years) occurred from time to time. A first check indicated that this noisy data occurred only during precipitation periods and temperatures slightly above 0 °C, possibly with snow grain. A separate meeting was arranged for 27 Feb 2018, discussing sonic data at CH-OE2.
Tech-SRP (TB): There is a problem with one of the HFP but only showing on some occasions. We assume due to a bad positioning, water accumulation on the field or erosion. Replacement of the SoilM profile is necessary. Change Sensors? Delta T PR2 à needs validation from a scientific side. Installed at FRU.


You are here: Short Report - QA/QC Meeting Feb 2018

Last Updated on 29 May 2021 19:36

Leave a Reply

You must be logged in to post a comment.